OpenStudy (anonymous):

Solve the problem. Use π = 3.14 when necessary. At $4.50 per m3, how much will it cost to fill an ice chest with dimensions of 3.2/3m x 7m x3.1/3m Round to the nearest cent. A. $385.00 B. $201.67 C. $350.00 D. $808.50

OpenStudy (pfenn1):

do you mean the dimensions are 3 2/3 m by 7 m by 3 1/3 m?

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

Is the ice chest rectangular or round? If it isn't round, why do we need pi?

OpenStudy (anonymous):

rectangular

OpenStudy (pfenn1):

The volume is given by Vol= height times width times depth. Multiply it out to get the total volume in cubic meters. Multiply the volume by the cost per cubic meters to get the total cost.

OpenStudy (anonymous):

7.71m^3 x 4.50m3 = 34.69m^3

OpenStudy (pfenn1):

Where did you get 7.71 cubic meters? my number was much larger than that.

OpenStudy (anonymous):

ya I got 383.80

OpenStudy (pfenn1):

you got 383.80 for what?

OpenStudy (anonymous):

(3.66 m) x (7 m) x (3.33 m) = 85.31 x 4.50m3

OpenStudy (pfenn1):

much better.

OpenStudy (anonymous):

and rounded it off 383.80 is that right

OpenStudy (pfenn1):

Well, I got 85.56 x 4.50 = 385 just some rounding errors. I multiplied \[\left( 11 \over 3 \right)\times7 \times \left( 10 \over 3 \right)=\left( 770\over 9 \right)\]

OpenStudy (pfenn1):

I don't mean rounding errors, I mean a difference because of rounding

OpenStudy (anonymous):

ok thank you for help me understand how to this.

OpenStudy (pfenn1):

you're welcome